number between 1-20 odds of getting an even number
Answer by jim_thompson5910(35256)

There are 20 numbers total and there are 10 even numbers in the range 1-20 (2,4,6,8,10,12,14,16,18,20)

P(Even) = # of evens/# total = 10/20 = 1/2


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim